OpenStudy (anonymous):

rewrite terms in sine and cosine only: sec X - cosX/(1+sin X)

OpenStudy (michele_laino):

we can substitute this: \[\sec x = \frac{1}{{\cos x}}\]

OpenStudy (michele_laino):

so we can write: \[\sec x - \frac{{\cos x}}{{1 + \sin x}} = \frac{1}{{\cos x}} - \frac{{\cos x}}{{1 + \sin x}}\]

OpenStudy (anonymous):

how would i simplify it?

OpenStudy (michele_laino):

hint: the common denominator is: \[\cos x\left( {1 + \sin x} \right)\]

OpenStudy (michele_laino):

what is: \[\frac{{\cos x\left( {1 + \sin x} \right)}}{{\cos x}} = ...?\]

OpenStudy (anonymous):

how is cos the denominator?

OpenStudy (michele_laino):

it is an intermediate step

OpenStudy (michele_laino):

the denominator of our new equivalent fraction is: \[{\cos x\left( {1 + \sin x} \right)}\]

OpenStudy (anonymous):

then i get \[\frac{1+\sin X-\cos ^{2}X\ }{ cosX(1+sinX) }\]

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

now, we can use this identity: \[{\left( {\cos x} \right)^2} = \left( {1 - \sin x} \right)\left( {1 + \sin x} \right)\]

OpenStudy (anonymous):

oh yay thank you so much !!!

OpenStudy (michele_laino):

:)